# Bistability and chaos in the discrete two-gene Andrecut-Kauffman model

### v1

@@ Abstract.
We conduct numerical analysis of the 2-dimensional discrete-time gene expression model originally introduced by Andrecut and Kauffman (Phys. Lett. A 367: 281–287, 2007). In contrast to the previous studies, we analyze the dynamics with different reaction rates $\alpha_{1}$ and $\alpha_{2}$ for each of the two genes under consideration. We explore bifurcation diagrams for the system with $\alpha_{1}$ varying in a wide range and $\alpha_{2}$ fixed. We detect chaotic dynamics by means of a positive maximum Lyapunov exponent and we scan through selected parameters to detect those combinations for which chaotic dynamics can be found in the system. Moreover, we find bistability in the model, that is, the existence of two disjoint attractors. Both situations are interesting from the point of view of applications, as they imply unpredictability of the system. Finally, we show some specific values of parameters of the model in which the two attractors are of different kind (a periodic orbit and a chaotic attractor) or of the same kind (two periodic orbits or two chaotic attractors).
@@
gene expression, discrete two-dimensional model, maximum Lyapunov exponent, chaos, bistability
@@
37N25, 37D45, 37G35, 65P20, 92C42
@@ 1. Introduction
Modeling gene expression plays a key role in understanding the regulatory mechanisms occurring within cells and the dynamics of biological processes at the molecular level. The development of mathematical models that accurately reflect the course of these processes allows for predicting the behavior of biological systems and contributes to the advancement of therapeutic strategies and genetic engineering. Gene expression models enable the study of complex networks of interactions among genes, transcripts, and proteins, which in turn contributes to a better understanding of such phenomena as cellular responses to stress and cell differentiation processes.
@@
Despite the undeniable complexity of the chemical reactions and kinetics involved in transcription and translation processes, the expression of a single gene does not reflect the intricacies of processes occurring within a living cell. When more genes are present in a system, their corresponding proteins can act as transcription factors binding to different promoters, creating an extensive network of dependencies. Due to their significance for all biological processes, numerous attempts have been made over the past decades to accurately model gene regulatory networks.
@@
The first approach to modeling gene expression involves continuous models, in which the system’s dynamics are described using systems of differential equations. This type of modeling allows for capturing the variables that describe the concentrations of gene products, such as mRNA or proteins, in a smooth manner over time. Among the classical models of this type are systems of differential equations that form the basis for analyzing dynamic dependencies among various components of a biological system, such as Goodwin’s model (see [Goo65]) and the Michaelis-Menten model (see [MM13]), which have been further developed by many authors. Such systems constitute effective tools for studying kinetic reactions and gene transcription, although their limitation is the necessity of assuming an immediate response of the system to changes in parameters. Therefore, in order to better capture biological realities, models incorporating time delays are considered. Delay models take into account time shifts that reflect the actual phenomena occurring during the synthesis of proteins and mRNA. This approach allows for a more precise reflection of processes where natural delays exist between the initiation of transcription and protein synthesis.
@@
The introduction of time delays into gene expression models has been proposed in numerous studies. For example, the authors of [BB12], [BB20] and [BB23] analyze continuous mathematical models of Hes1 gene expression, incorporating various aspects such as time delays, feedback mechanisms, and dimerization processes. The paper [BB12] introduces separate delays for transcription and translation, highlighting the impact of the total delay on stability and the occurrence of Hopf bifurcation. In [BB20], the model is expanded to include different time scales and is reduced to the classical Hes1 model, examining equilibrium stability and dynamics. Finally, further development of the model is proposed in [BB23] by including dimerization, and analysis of the impact of delays on bifurcations and stability of limit cycles is conducted, providing deeper insights into oscillations in Hes1 gene expression. Similarly, the authors of [PBB18] investigate a generalized p53-Mdm2 model and demonstrate that oscillations in gene expression can arise with or without time delays, depending on specific stability conditions and bifurcation criteria.
@@
The groundwork for the model analyzed in our research was laid out by Kauffman in 1969 in the form of boolean networks [Kau69]. A boolean network is a finite set of boolean variables changing in discrete time steps according to predetermined rules. A single network is able to possess several distinct modes of behavior, analogous to one genome creating specialized cell types. Further analysis of random boolean networks was facilitated in 2005 when Andrecut used mean field theory to calculate probabilities of node states in a network given only global parameters. Moreover, he has calculated Lyapunov exponents of such networks as a function of connectivity, and has shown that for the purpose of investigating chaotic behavior the discrete network can be replaced with a continuous map. This allowed for application of Sharkovskii’s theorem to find possible orbits of the system [And05].
@@
In 2006, Andrecut and Kauffman proposed a continuous model in which individual reactions between genes have been replaced with their average results, reducing computational complexity and making the outcome deterministic, as opposed to the more realistic stochastic simulations. The regulatory network is represented as a system of differential equations, with exact formulae derived directly from the underlying chemical reactions. The model accounts for $N$ genes and protein multimers up to length $n$, including mixed protein chains. Fourier analysis of numerical solutions has shown that this model admits a range of different behaviors, such as stationary solutions, periodic orbits, and chaotic behavior as well [AK06].
@@
A more approachable, discrete-time model was proposed a year later in [AK07]. The number of genes was fixed at $2$ and only homo-multimers of length exactly $n$ were taken into consideration. This led to a system of two mutually dependent equations. Despite the considerable simplification, the new model seems to represent the complex dynamics of the original system well, displaying different modes of behavior under various conditions. Combined with the relative ease of simulating a discrete model, this has encouraged further research.
@@
In our work, we focus on the discrete Andrecut-Kauffman model. However, we consider a generalized form in which a different reaction rate may be associated to each gene (see Section 3). This distinction allows for a more precise representation of gene expression dynamics. Specifically, instead of a single, general rate $\alpha$, inclusion of $\alpha_{1}$ and $\alpha_{2}$ in the equations of the model enables capturing various regulatory mechanisms that may influence transcription processes. Additionally, we incorporate the Lyapunov exponent into our analysis, which enables us to measures the system’s stability and sensitivity to initial conditions, an essential factor for understanding the complex dynamics of discrete models in a biological context.
@@
The purpose of our research is to investigate the dynamics encountered in the generalized discrete Andrecut-Kauffman model important for applications. We put emphasis on the existence of chaotic dynamics, and also on the phenomenon of bistability, that is, the existence of two distinct attractors, possibly of different types, such as an attracting periodic orbit and a chaotic attractor. Both phenomena are of undeniable importance from the biological point of view. Chaotic dynamics implies unpredictability of the long-term behavior of the system. Bistability implies the fact that the system can stabilize in two different ways, depending on the initial conditions.
@@
The paper is organized as follows. In Section 2, we explain the gene expression mechanisms that yield the difference equations that we analyze. The equations are introduced in Section 3. Then in Section 4, we establish an absorbing set that captures all the bounded dynamics. In Section 5, we show evidence that admitting different reaction rates proposed in Section 3.3 may result in qualitative change of dynamics and is thus an important direction of investigation. In Section 6, we show and analyze bifurcation diagrams of the system with one of the reaction rates varying in a wide interval while keeping the other rate fixed. Then we introduce a numerical procedure for the computation of the maximum Lyapunov exponent in our $2$-dimensional model in Section 7, and we apply it to search for chaotic dynamics in Section 8. We then analyze the maximum Lyapunov exponents for a wide range of the reaction rates in Section 9. Then in Section 10, we discuss the asymmetry found in the computation of the maximum Lyapunov exponent, which yields to the detection of three kinds of bistability in the system that we discuss in Section 11. The paper concludes with a discussion in Section 12, where we address the significance of our findings for understanding gene expression dynamics, emphasizing the role of parameters in system stability and its potential to exhibit chaotic behavior. We also suggest future research directions, such as the analysis of bifurcations, unstable fixed points, and global dynamics in the context of systems biology and controlled genetic systems.
@@ 2. Gene expression mechanisms
Gene expression encompasses all the processes that determine the amount of a specific mRNA and protein produced within a cell. This series of processes includes two fundamental stages: transcription, where a gene’s DNA sequence is transcribed into mRNA, and translation, where the mRNA is subsequently translated into a protein. Gene expression often serves as a foundational aspect in cell biology research, providing insight into mechanisms at both microscopic and molecular scales. Understanding how gene expression is regulated and coordinated is crucial for grasping the complexity of cellular systems [LSZ09].
@@ 2.1. Transcription
Transcription is the process where enzymes use one strand of DNA within a gene as a template to produce messenger RNA (mRNA). RNA polymerase, aided by proteins called transcription factors, binds to a specific sequence within the gene known as the promoter and separates the two DNA strands. The template strand, or antisense strand, is used to generate the mRNA, while the other strand is the nontemplate or sense strand. RNA polymerase initiates mRNA synthesis at the start codon without needing a primer and moves downstream along the gene in a process called elongation. It synthesizes the mRNA by reading the antisense strand and generating the mRNA, adding RNA nucleotides as it goes. This process is similar to how DNA polymerase synthesizes DNA, except RNA is being synthesized with ribose instead of deoxyribose and uracil instead of thymine. Unlike DNA replication, RNA polymerase re-zips the DNA behind it, keeping only 10 to 20 bases exposed at a time. When RNA polymerase reaches the end of the gene, termination occurs, the enzyme detaches, and the DNA returns to its original state. The resulting mRNA, after a few modifications during RNA processing, leaves the nucleus and moves into the cytoplasm, where it finds a ribosome [AJL+02].
@@
Figure 1. Schematic representation of the DNA transcription.
@@ 2.2. Translation
Translation occurs in the ribosome, where the mRNA serves as a code for a specific protein. Each set of three bases on the mRNA, called codons, codes for a specific anticodon carried by a transfer RNA (tRNA), each linked to a particular amino acid. The arrangement of nucleotides into codons is called the reading frame. With four bases and codons made of three bases each, there are 64 possible codons, enough to code for all necessary amino acids. Each codon corresponds to a particular amino acid; AUG is the start codon, initiating translation by coding for methionine, and there are three stop codons (UAG, UAA, UGA) that terminate translation.
During translation, the small ribosomal subunit binds to the mRNA and an initiator tRNA, which adheres to the start codon. The large ribosomal subunit then joins to complete the translation initiation complex. The tRNA corresponding to the next codon enters the ribosome, carrying an amino acid that binds covalently to the methionine from the initiator tRNA. The first tRNA detaches and leaves the ribosome, which shifts to accommodate the next tRNA. This process continues along the mRNA, with tRNAs entering and exiting the ribosome according to the codons on the mRNA, and the polypeptide chain grows. When a stop codon is reached, the completed polypeptide is released, likely entering a cell organelle for folding and further modification [AJL+02].
@@
Through this two-step process, DNA is transcribed into mRNA, and then this mRNA is translated into a protein. Since every gene codes for a specific protein and proteins constitute much of an organism’s structure and function, from muscle and organ tissue to receptors and enzymes, this is how DNA carries the code for a living organism.
@@ 3. The Andrecut-Kauffman model
In molecular biology, the flow of genetic information follows this pathway of transcription and translation, where DNA is first transcribed into messenger RNA (mRNA), which is subsequently translated into protein. Each gene within DNA encodes a specific protein, essential for an organism’s structure and function, forming cells, tissues, and organs, and playing critical roles in processes like muscle structure and enzymatic activity. Thus, DNA serves as the blueprint for life by directing the synthesis of the proteins required for a functioning organism.
@@
Andrecut and Kauffman [AK07] introduced a simplified model of gene expression that merges the processes of transcription and translation described in Section 2 into a single reaction. This approach simplifies the complex biological processes by treating them as one integrated step, making it easier to analyze the system’s dynamics.
The main motivation for the introduction of the discrete Andrecut-Kauffman model is the set of chemical reactions for the two-gene system, which covers the complex steps of gene expression. In this model, the key reactions include:
@@
    captures how multimeric complexes ($M_{n}$) bind to promoter sites ($P$), thereby blocking further transcription. Such promoter blocking mechanisms are essential for regulatory feedback, where the complex ($P(M_{n})$) can inhibit continued gene expression.
@@
    describes the degradation of the protein monomers ($M$), a vital aspect of gene regulation that affects the level of gene expression by controlling the lifespan of protein molecules.
@@
In this framework, the dynamic relationships between the expression levels of two genes, labeled as $x$ and $y$, are explored through their interactions and feedback mechanisms. The model defines these relationships with the following equations:
@@
Here, $x_{t}$ and $y_{t}$ represent the concentration levels of transcription factor proteins for the genes $x$ and $y$ at time $t$, respectively. The parameters represent properties of the chemical reactions discussed in Section 3.1. The parameter $\alpha=\Delta t\cdot k$ reflects the rate of combined transcription and translation over a fixed time interval. The parameter $\beta=1-\Delta t\cdot k^{\delta}$ depends on the rate of protein degradation for a fixed time interval. Finally, $\varepsilon$ is a parameter that describes the coupling between the genes under the assumption that it is symmetric [AK06]. The exponent $n$ denotes the number of monomers of a given protein that are subject to the multimerization reaction. These parameters were considered in [AK07] in the following ranges: $n\in\{1,2,3,4\}$, $\alpha\in[0,100]$, $\beta\in[0,1)$, and $\varepsilon\in[0,1]$.
@@
The model proposed in [AK07] assumes many simplifications, including the assumption that the genes are identical both in terms of rate of the gene expression reaction, and the degradation of protein monomers as well. In [SS19, SKR+23], the Andrecut-Kauffman model was considered with different rates of degradation $\beta$ for each gene. In our research, however, we will deal with different reaction rates $\alpha$ for each gene instead. In other words, we use $\alpha_{1}$ and $\alpha_{2}$ in the two equations instead of having the common rate $\alpha$. We thus consider the following model:
@@
Moreover, the definition of the parameter $\alpha=\Delta t\cdot k$ in no way limits its value to the range $[0,100]$ considered previously; therefore, in what follows, we consider larger ranges of the parameters $\alpha_{1}$ and $\alpha_{2}$ in order to examine the dynamics of the system more broadly.
In the original analysis, Andrecut and Kauffman [AK07] indicate that for $n=1$ and $n=2$ the system exhibits no chaotic behavior and only produces stable points or cycles. We have conducted low resolution simulations to validate these findings using methods described in Section 7. Indeed, computed Lyapunov exponents were negative in the considered parameter ranges. Therefore, we decided to restrict our attention to $n=3$ and $n=4$ in our research.
4. An absorbing set for the system
Before conducting numerical investigation of the system (2), it is worth to know that within the regime of parameters of our interest and among all the relevant initial conditions there are no trajectories that become unbounded in forward time. Moreover, we would like to know that all the attractors and other sets that represent long-term dynamics are located within a specific bounded region of the phase space, so that we can conduct informed numerical simulations and not worry that we miss any important part of the dynamics. The following notion is useful for this purpose.
Definition 1 (Absorbing set, see [PG24, Definition 1]).
A set $P\subset X$ is called an absorbing set for a dynamical system $f\colon X\to X$ if for every $x\in X$ there exists $N_{0}>0$ such that $f^{n}(x)\in P$ for all $n\geq N_{0}$.
Note that although every forward trajectory in the system enters an absorbing set in a finite number of steps, the actual number of the steps may depend on the initial conditions. Moreover, every absorbing set contains all the attractors of the system (both local and global). Even more, it contains all the recurrent dynamics present in the system, such as stable and unstable fixed points and periodic orbits.
Note that the entire phase space satisfies the definition of an absorbing set for a dynamical system. However, we are interested in finding a possibly small bounded absorbing set in order to be able to restrict our numerical simulations to a manageable subset of the phase space and be sure we obtain reliable results. The following proposition defines a useful absorbing set for the model (2).
Proposition 2.
Let $c_{1},c_{2}>0$ be arbitrary numbers (in particular, they can be very small). Define $b_{i}:=(\alpha_{i}+c_{i})/(1-\beta)$ for $i\in\{1,2\}$. Then the set $R_{c}:=[0,b_{1}]\times[0,b_{2}]$ is an absorbing set for the model (2) with the parameters $\alpha_{1},\alpha_{2}\geq 0$, $\beta\in[0,1)$, and any $\varepsilon\in[0,1]$ and $n\in\mathbb{N}$. Moreover, the set $R_{c}$ is positively invariant.

### v2

@@ Abstract.
We conduct numerical analysis of the 2-dimensional discrete-time gene expression model originally introduced by Andrecut and Kauffman (Phys. Lett. A 367: 281–287, 2007). In contrast to the previous studies, we analyze the dynamics with different reaction rates $\alpha_{1}$ and $\alpha_{2}$ for each of the two genes under consideration. We explore bifurcation diagrams for the model with $\alpha_{1}$ varying in a wide range and $\alpha_{2}$ fixed. We detect chaotic dynamics by means of the positive maximum Lyapunov exponent and we scan through selected parameters to detect those combinations for which chaotic dynamics can be found in the model. Moreover, we find bistability in the model, that is, the existence of two disjoint attractors. Both situations are interesting from the point of view of applications, as they imply unpredictability of the dynamics encountered. Finally, we show some specific values of parameters of the model in which the two attractors are of different kind (a periodic orbit and a chaotic attractor) or of the same kind (two periodic orbits or two chaotic attractors).
@@
Gene expression, discrete two-dimensional model, maximum Lyapunov exponent, chaos, bistability.
@@
Primary: 37N25, 65P20; Secondary: 37D45, 37G35, 92C42.
Both authors contributed equally to the research
Corresponding author, https://pg.edu.pl/p/agnieszka-bartlomiejczyk-13028
This is the author accepted manuscript (postprint) of the paper published in Discrete and Continuous Dynamical Systems – Series B (DCDS-B), Vol. 30, No. 11, November 2025, pp. 4442–4461, https://doi.org/10.3934/dcdsb.2025028.
This postprint is licensed under the CC-BY-NC-ND 4.0 Creative Commons License.
Mikołaj Rosman†1 Michał Palczewski†1,2 Paweł Pilarczyk1,3 and Agnieszka Bartłomiejczyk∗1,4
1Faculty of Applied Physics and Mathematics, Gdańsk University of Technology, ul. Narutowicza 11/12, 80-233 Gdańsk, Poland
2Doctoral School, Gdańsk University of Technology, ul. Narutowicza 11/12, 80-233 Gdańsk, Poland
3Digital Technologies Center, Gdańsk University of Technology, ul. Narutowicza 11/12, 80-233 Gdańsk, Poland
4BioTechMed Center, Gdańsk University of Technology, ul. Narutowicza 11/12, 80-233 Gdańsk, Poland
@@ 1. Introduction
Modeling gene expression plays a key role in understanding the regulatory mechanisms within cells and the dynamics of biological processes at the molecular level. The development of mathematical models that accurately reflect the course of these processes allows for predicting the behavior of biological systems and contributes to the advancement of therapeutic strategies and genetic engineering. Gene expression models enable the study of complex networks of interactions between genes, transcripts, and proteins, which in turn contributes to a better understanding of phenomena such as cellular responses to stress and cell differentiation processes.
@@
Despite the undeniable complexity of the chemical reactions and kinetics involved in transcription and translation processes, the expression of a single gene does not reflect the intricacies of processes occurring within a living cell. When more genes are present in a system, their corresponding proteins can act as transcription factors, binding to different promoters, creating an extensive network of dependencies. Due to their significance for all biological processes, numerous attempts have been made over the past decades to accurately model gene regulatory networks.
@@
The first approach to modeling gene expression involves continuous-time models, in which the dynamics is described using systems of differential equations. This type of modeling allows for capturing the variables that describe the concentrations of gene products, such as mRNA or proteins, in a smooth manner over time. Among the classical models of this type are systems of differential equations. These systems form the basis for analyzing dynamic dependencies among various components of a biological system, such as Goodwin’s model (see [10]) and the Michaelis-Menten model (see [16]), which have been further developed by many authors. Such models constitute effective tools for studying kinetic reactions and gene transcription, although their limitation is the necessity of assuming an immediate response of the system to changes in parameters. Therefore, in order to better capture biological realities, models that incorporate time delays are considered. Delay models take into account time shifts that reflect the actual phenomena taking place during the synthesis of proteins and mRNA. This approach allows for a more precise reflection of processes where natural delays exist between the initiation of transcription and protein synthesis.
@@
The introduction of time delays into gene expression models has been proposed in numerous studies. For example, the authors of [7, 8, 9] analyze continuous mathematical models of Hes1 gene expression, incorporating various aspects such as time delays, feedback mechanisms, and dimerization processes. The paper [9] introduces separate delays for transcription and translation, highlighting the impact of the total delay on stability and the occurrence of a Hopf bifurcation. In [7], the model is expanded to include different time scales and is reduced to the classical Hes1 model, examining equilibrium stability and dynamics. Finally, further development of the model is proposed in [8] by including dimerization, and analysis of the impact of delays on bifurcations and stability of limit cycles is being conducted, which results in deeper insight into oscillations in Hes1 gene expression. Similarly, the authors of [19] investigate a generalized p53-Mdm2 model and demonstrate that oscillations in gene expression can arise with or without time delays, depending on specific stability conditions and bifurcation criteria.
@@
The groundwork for the model analyzed in our research was laid out by Kauffman in 1969 in the form of Boolean networks [11]. A Boolean network is a finite set of Boolean variables changing in discrete time steps according to predetermined rules. A single network can have several distinct modes of behavior, by analogy to one genome creating specialized cell types. Further analysis of random Boolean networks was facilitated in 2005 when Andrecut used mean field theory to calculate probabilities of node states in a network given only global parameters. Moreover, he has calculated Lyapunov exponents of such networks as a function of connectivity, and has shown that for the purpose of investigating chaotic behavior the discrete network can be replaced with a continuous map. This allowed for application of Sharkovskii’s theorem to find possible orbits present in the model [3].
@@
In 2006, Andrecut and Kauffman proposed a continuous model in which individual reactions between genes have been replaced with their average results, reducing computational complexity and making the outcome deterministic, as opposed to the more realistic stochastic simulations. The regulatory network is represented as a system of differential equations, with exact formulae derived directly from the underlying chemical reactions. The model accounts for $N$ genes and protein multimers up to length $n$, including mixed protein chains. Fourier analysis of numerical solutions has shown that this model admits a range of different behaviors, such as stationary solutions, periodic orbits, and chaotic behavior as well [4].
@@
A more approachable, discrete-time model was proposed a year later in [5]. The number of genes was fixed at $2$ and only homo-multimers of length exactly $n$ were taken into consideration. This led to a system of two mutually dependent equations. Despite the considerable simplification, the new model seems to represent the complex dynamics of the original model well, displaying different modes of behavior under various conditions. Combined with the relative ease of simulating a discrete model, this has encouraged further research.
@@
In our work, we focus on the discrete Andrecut-Kauffman model. However, we consider a generalized form in which a different reaction rate may be associated to each gene (see Section 3). This distinction allows for a more precise representation of gene expression dynamics. Specifically, instead of a single, general rate $\alpha$, inclusion of $\alpha_{1}$ and $\alpha_{2}$ in the equations of the model enables capturing various regulatory mechanisms that may influence transcription processes. Additionally, we incorporate the Lyapunov exponent into our analysis, which enables us to measure the model’s stability and sensitivity to initial conditions, an essential factor for understanding the complex dynamics of discrete models in a biological context.
@@
The purpose of our research is to investigate the dynamics encountered in the generalized discrete Andrecut-Kauffman model important for applications. We put emphasis on the existence of chaotic dynamics, and also on the phenomenon of bistability, that is, the existence of two distinct attractors, possibly of different types, such as an attracting periodic orbit and a chaotic attractor. Both phenomena are of undeniable importance from the biological point of view. Chaotic dynamics implies unpredictability of the long-term behavior of the modeled system. Bistability implies the fact that the system can stabilize in two different ways depending on the initial conditions.
@@
The paper is organized as follows. In Section 2, we explain the gene expression mechanisms that yield the difference equations that we analyze. The equations are introduced in Section 3. Then in Section 4, we establish an absorbing set that captures all the bounded dynamics. In Section 5, we show evidence that admitting different reaction rates proposed in Section 3.3 can result in qualitative change of dynamics and is thus an important direction of investigation. In Section 6, we show and analyze bifurcation diagrams of the model with one of the reaction rates varying in a wide range while keeping the other rate fixed. Then we introduce a numerical procedure for the computation of the maximum Lyapunov exponent in our $2$-dimensional model in Section 7, and we apply it to the search for chaotic dynamics in Section 8. We then analyze the maximum Lyapunov exponents for a wide range of reaction rates in Section 9. Then in Section 10, we discuss the asymmetry found in the computation of the maximum Lyapunov exponent, which yields to the detection of three kinds of bistability in the model that we discuss in Section 11. The paper concludes with a discussion in Section 12, where we address the importance of our findings for understanding gene expression dynamics, emphasizing the role of parameters in the stability of the analyzed model and its potential to exhibit chaotic behavior. We also suggest future research directions, such as the analysis of bifurcations, unstable fixed points, and global dynamics in the context of Systems biology and controlled genetic systems.
@@ 2. Gene expression mechanisms
Gene expression encompasses all the processes that determine the amount of a specific mRNA and protein produced within a cell. This series of processes includes two fundamental stages: transcription, where a gene’s DNA sequence is transcribed into mRNA, and translation, where the mRNA is subsequently translated into a protein. Gene expression often serves as a foundational aspect in cell biology research, providing insight into mechanisms both at the microscopic and molecular scale. Understanding how gene expression is regulated and coordinated is crucial for grasping the complexity of cellular systems [12].
@@ 2.1. Transcription
Transcription is the process in which enzymes use one strand of DNA within a gene as a template to produce messenger RNA (mRNA). RNA polymerase, aided by proteins called transcription factors, binds to a specific sequence within the gene known as the promoter and separates the two DNA strands. The template strand, or the antisense strand, is used to generate the mRNA, while the other strand is the nontemplate or sense strand. RNA polymerase initiates mRNA synthesis at the start codon without needing a primer and moves downstream along the gene in a process called elongation. It synthesizes the mRNA by reading the antisense strand and generating the mRNA, adding RNA nucleotides as it goes. This process is similar to how DNA polymerase synthesizes DNA, except that RNA is being synthesized with ribose instead of deoxyribose and uracil instead of thymine. Unlike DNA replication, RNA polymerase re-zips the DNA behind it, keeping only 10 to 20 bases exposed at a time. When RNA polymerase reaches the end of the gene, termination occurs, the enzyme detaches, and the DNA returns to its original state. The resulting mRNA, after a few modifications during RNA processing, leaves the nucleus and moves to the cytoplasm, where it finds a ribosome [1].
@@
Figure 1. Schematic representation of DNA transcription.
@@ 2.2. Translation
Translation occurs in the ribosome, where mRNA serves as a code for a specific protein. Each set of three bases on the mRNA, called codons, encodes a specific anticodon carried by a transfer RNA (tRNA), each linked to a particular amino acid. The arrangement of nucleotides into codons is called the reading frame. With four bases to choose, and codons consisting of three bases each, there are $64$ possible codons, enough to encode all the necessary amino acids. Each codon corresponds to a particular amino acid; AUG is the start codon, initiating translation by coding for methionine, and there are three stop codons (UAG, UAA, UGA) that terminate translation.
During translation, the small ribosomal subunit binds to the mRNA and an initiator tRNA, which adheres to the start codon. The large ribosomal subunit then joins to complete the translation initiation complex. The tRNA corresponding to the next codon enters the ribosome, carrying an amino acid that binds covalently to the methionine from the initiator tRNA. The first tRNA detaches and leaves the ribosome, which shifts to accommodate the next tRNA. This process continues along the mRNA, with tRNAs entering and exiting the ribosome according to the codons on the mRNA, and the polypeptide chain grows. When a stop codon is reached, the completed polypeptide is released, likely entering a cell organelle for folding and further modification [1].
@@
Through this two-step process, DNA is transcribed into mRNA, and then this mRNA is translated into a protein. Since every gene encodes a specific protein and proteins constitute much of an organism’s structure and function, from muscle and organ tissue to receptors and enzymes, this is how DNA carries the code for a living organism.
@@ 3. The Andrecut-Kauffman model
In molecular biology, the flow of genetic information follows this transcription and translation pathway, where DNA is first transcribed into messenger RNA (mRNA), which is subsequently translated into protein. Each gene within DNA encodes a specific protein, essential for an organism’s structure and function, forming cells, tissues, and organs, and playing critical roles in processes like muscle structure and enzymatic activity. Thus, DNA serves as the blueprint for life by directing the synthesis of the proteins required for a functioning organism.
@@
Andrecut and Kauffman [5] introduced a simplified model of gene expression that merges the processes of transcription and translation described in Section 2 into a single reaction. This approach simplifies the complex biological processes by treating them as one integrated step, making it easier to analyze the model’s dynamics.
The main motivation for the introduction of the discrete Andrecut-Kauffman model is the set of chemical reactions for the two-gene system, which covers the complex steps of gene expression. In this model, the key reactions include the following.
@@
    captures how multimeric complexes ($M_{n}$) bind to promoter sites ($P$), thus blocking further transcription. Such promoter blocking mechanisms are essential for regulatory feedback, where the complex ($P(M_{n})$) can inhibit continued gene expression.
@@
    describes the degradation of protein monomers ($M$), a vital aspect of gene regulation that affects the level of gene expression by controlling the lifespan of protein molecules.
@@
In this framework, the dynamic relationships between the expression levels of two genes, labeled $x$ and $y$, are explored through their interactions and feedback mechanisms. The model defines these relationships with the following equations:
@@
Here, $x_{t}$ and $y_{t}$ represent the concentration levels of transcription factor proteins for the genes $x$ and $y$ at time $t$, respectively. The parameters represent properties of the chemical reactions discussed in Section 3.1. The parameter $\alpha=\Delta t\cdot k$ reflects the rate of combined transcription and translation over a fixed time interval. The parameter $\beta=1-\Delta t\cdot k^{\delta}$ depends on the rate of protein degradation for a fixed time interval. Finally, $\varepsilon$ is a parameter that describes the coupling between the genes under the assumption that it is symmetric [4]. The exponent $n$ denotes the number of monomers of a given protein that are subject to the multimerization reaction. These parameters were considered in [5] in the following ranges: $n\in\{1,2,3,4\}$, $\alpha\in[0,100]$, $\beta\in[0,1)$, and $\varepsilon\in[0,1]$.
@@
The model proposed in [5] assumes many simplifications, including the assumption that the genes are identical in terms of the rate of gene expression reaction, and the degradation of protein monomers as well. In [20, 22], the Andrecut-Kauffman model was considered with different rates of degradation $\beta$ for each gene. However, in our research, we will instead deal with different reaction rates $\alpha$ for each gene. In other words, we use $\alpha_{1}$ and $\alpha_{2}$ in the two equations instead of having the common rate $\alpha$. We thus consider the following model:
@@
Moreover, the definition of the parameter $\alpha=\Delta t\cdot k$ in no way limits its value to the range $[0,100]$ considered previously; therefore, in what follows, we consider larger ranges of the parameters $\alpha_{1}$ and $\alpha_{2}$ in order to examine the dynamics of the model more broadly.
In the original analysis, Andrecut and Kauffman [5] indicate that for $n=1$ and $n=2$ the model exhibits no chaotic behavior and only produces stable points or cycles. We have conducted low resolution simulations to validate these findings using methods described in Section 7. Indeed, the computed Lyapunov exponents were negative in the range of parameters considered. Therefore, we decided to limit our attention to $n=3$ and $n=4$ in our research.
4. An absorbing set for the model
Before conducting numerical investigation of model (2), it is worth to know that within the regime of parameters of our interest and among all the relevant initial conditions there are no trajectories that become unbounded in forward time. Moreover, we would like to know that all the attractors and other sets that represent long-term dynamics are located within a specific bounded region of the phase space, so that we can conduct informed numerical simulations and not worry that we miss any important part of the dynamics. The following notion is useful for this purpose.
Definition 4.1 (Absorbing set, see [17, Definition 1]).
A set $P\subset X$ is called an absorbing set for a dynamical system generated by $f\colon X\to X$ if for every $x\in X$ there exists $N_{0}>0$ such that $f^{n}(x)\in P$ for all $n\geq N_{0}$.
Note that although every forward trajectory in the dynamical system enters an absorbing set in a finite number of steps, the actual number of steps may depend on the initial conditions. Moreover, every absorbing set contains all the attractors of the system (both local and global). Moreover, it contains all the recurrent dynamics present in the system, such as stable and unstable fixed points and periodic orbits.
Note that the entire phase space satisfies the definition of an absorbing set for a dynamical system. However, we are interested in finding a possibly small bounded absorbing set in order to be able to restrict our numerical simulations to a manageable subset of the phase space and still ensure that we obtain reliable results. The following proposition defines a useful absorbing set for model (2).
Proposition 4.2.
Let $c_{1},c_{2}>0$ be arbitrary numbers (in particular, they can be very small). Define $b_{i}:=(\alpha_{i}+c_{i})/(1-\beta)$ for $i\in\{1,2\}$. Then the set $R_{c}:=[0,b_{1}]\times[0,b_{2}]$ is an absorbing set for model (2) with $\alpha_{1},\alpha_{2}\geq 0$, $\beta\in[0,1)$, and any $\varepsilon\in[0,1]$ and $n\in\mathbb{N}$. Moreover, the set $R_{c}$ is positively invariant.